OpenStudy (anonymous):

What angle (in degrees) is equivalent to 3p/ 2? p/ 4?

OpenStudy (anonymous):

\[degrees = radians \times \frac{ 180 }{ \pi }\]

OpenStudy (anonymous):

Whats the asnwer.?

OpenStudy (anonymous):

That's for you to solve. Just plug in the value in "radians"...

OpenStudy (anonymous):

whatt if i just put Pi=180? will it be okay?

OpenStudy (anonymous):

No you want to cancel out the radians so you are left with degrees.